OpenStudy (anonymous):

Which expressions could complete this equation so that it has one solution? Choose exactly two answers that are correct.

OpenStudy (anonymous):

im here

OpenStudy (anonymous):

whats the equation

OpenStudy (anonymous):

The equation is 4(4x-3)-6x=?

OpenStudy (anonymous):

A. 12X-12-2X B. 9X-12 C. 4X+2(3X-6) D. 4(2X+3)-24

OpenStudy (welshfella):

try each one in turn for first one 4(4x-3)-6x = 12x - 12 - 2x 16x - 12 - 6x = 10x - 12 10x - 12 = 10x - 12 for which there are infinite solutions thats not the one

OpenStudy (welshfella):

B 10x - 12 = 9x - 12

OpenStudy (welshfella):

has this one got one solution?

OpenStudy (welshfella):

try to solve it first subtract 9x from both sides then add 12 to both sides

OpenStudy (anonymous):

okay so right now i have 9x-10=12+12

OpenStudy (anonymous):

and after that i got X=-24?